3)Let x=n+ε, where n is an integer and 0≤ε<1. If ⌊55x/7⌋=8n, then which of the following cannot be a value of x:

a) 14.3 

b) 37.7

c) 46.9

d) 51.6

4) Let x=n+ε, where n is an integer and 0≤ε<1. If ⌈33x/8⌉=5n, then which of the following is a valid value of x:

a) 5.8 

b) 4.7

c) 6.1

d) 2.5


1
Expert's answer
2021-05-07T09:25:57-0400

(i) As "x=n+\u03b5, 0\u2264\u03b5<1"


Given relation is- "\dfrac{55x}{7}=8n"

(a) Let x=14.3, then n=14 as "x=n+\u03b5"


Taking LHS"=\dfrac{55\times 14.3}{7}=112.35"


Taking RHS="8(14)=112"

(b) Let x=37.7, then n=37


Taking LHS="\dfrac{55\times 37.7}{7}=296.4"


Taking RHS="8(37)=296"


(c) Let x=46.9, then n=46


Taking LHS="\dfrac{55\times 46.9}{7}=368.5"


Taking RHS="8(46)=368"

(d) Let x=51.6, then n=51


Taking LHS="\dfrac{55\times 51.6}{7}=405.42"


Taking RHS="8(51)=408"


The difference between the LHS and RHS is more in part D,

Hence The value of x can not be 51.6


(ii) Given Relation is "\dfrac{33x}{8}=5n"


(a) Let x=5.8 , then n=5


Taking LHS ="\dfrac{33\times 5.8}{8}=23.925"


Taking RHS="5(5)=25"


(b) Let x=4.7 , then n=4


Taking LHS =\dfrac{33\times 4.7}{8}=19.3875

Taking RHS=5(4)=20


(c) Let x=6.1 , then n=6

Taking LHS "=\dfrac{33\times 6.1}{8}=25.1625"


Taking RHS="5(6)=30"


(d) Let x=2.5 , then n=2


Taking LHS ="\dfrac{33\times 2.5}{8}=10.3125"


Taking RHS="5(2)=10"


The difference between the RHS and LHS is significant in part (a) and (C)

Therefore x can not have the values 5.8 and 6.1